47. On his way to work, Alan passes an intersection with a traffic light that hasa 50% chance of being green for the direction in which he is travelling. a) What is the probability that this traffic light will be green as Alan reaches the intersection at least once a week?
Show work and a) 0.97 b ) 1 <— answers
b) What is the expected number of days before the light is green when Alan reaches the intersection?
Answer #1

P(X >= 1) = 1 - P(X =0)
= 1 - (1 - 0.5)^5    { 5 because there are 5 working days in a week}

= 0.96875